# BOP-D: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

### v1

@@ 5.2 Qualitative validation of the method
SpyroPose [7] RGB 61.2 75.3 68.3 1 57.8 71.6 64.7 1
LiePose Diffusion [17] RGB 60.1 92.2 76.1 2 48.3 76.1 62.2 2
T-LESS [11] Challenge: pose estimation of unseen objects
foundationpose [43] RGB-D 62.9 64.1 63.5 1 56.7 58.0 57.4 1
foundpose [34] RGB 55.0 62.3 58.6 2 45.5 52.1 48.8 5 ( $\downarrow$3)
gigaposemegapose [32, 22] RGB 54.3 62.0 58.1 3 45.3 52.3 48.8 4 ( $\downarrow$1)
gigaposegenflowkabsch [32, 29] RGB-D 56.3 58.6 57.5 4 49.4 51.7 50.6 2 ( $\uparrow$2)
gigaposemegapose [32] RGB 51.5 59.2 55.3 5 38.5 45.1 41.8 11 ( $\downarrow$6)
genflow-multihypo [29] RGB 50.9 57.9 54.4 6 45.9 52.4 49.1 3 ( $\uparrow$3)
genflow-multihypo [29] RGB 50.4 57.3 53.9 7 44.9 51.4 48.1 6 ( $\uparrow$1)
genflow-multihypo16 [29] RGB-D 52.8 54.8 53.8 8 46.7 48.6 47.7 7 ( $\uparrow$1)
sam6d [23] RGB-D 52.9 54.5 53.7 9 39.7 41.2 40.4 15 ( $\downarrow$6)
sam6d [23] RGB-D 52.1 53.5 5208 10 38.3 39.7 39.0 18 ( $\downarrow$8)
genflow-multihypo [29] RGB-D 51.6 53.6 52.6 11 45.2 47.2 46.2 8 ( $\uparrow$3)
foundpose [43] RGB 49.1 55.6 52.4 12 37.7 43.4 40.6 13 ( $\downarrow$1)
foundpose [43] RGB 49.1 55.6 52.3 13 37.7 43.4 40.5 14 ( $\downarrow$1)
cnos-fastsam10 [31] RGB 48.5 55.9 52.2 14 40.6 47.1 43.8 10 ( $\uparrow$4)
Table 1: Results of BOP 2023 on T-LESS [11]. We report the top contenders results of the pose estimation challenge (2023 edition) evaluated with $\mathbf{MSSD}$ and $\mathbf{MSPD}$ on both the official ground truth and our ground truth. We rank the methods with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ (similar to BOP, except we exclude $\mathbf{VSD}$ as BOP starts to abandon it) and re-rank them with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ computed on our ground truth distributions. This produces drastic changes in the rankings. We also show the same scores for the best mode of distribution estimation methods, however, both of them use ground truth bounding boxes instead of a detector, so direct comparison would be unfair. Finally, we apply the same study to the pose estimation of unseen objects challenge. Similarly, the rankings change. This highlights the importance to consider image level symmetries as we proposed, to evaluate accurately pose estimation methods. (names of the methods are automatically processed from BOP csv result files)
@@ 5.3 Single Pose Estimation Evaluation
Since May 2023, raw results of BOP Challenge [15] public submissions have become publicly available to allow in-depth analysis. In this experiment, we use them to compare the performance evaluations of state-of-the-art methods on T-LESS for both the original 6D annotations and our annotations. We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks. We also included two methods for the multi-modal pose distribution estimation task [7, 17] for which we kept only the mode with highest confidence mode. We consider their full predicted distributions in the next subsection. Since these methods are not part of BOP challenge, we retrained them on the T-LESS train set including both photo-realistic synthetic images and real images.
@@
We evaluate the current (July 2024) top contenders of the BOP Challenge on the T-LESS dataset [11] on the pose estimation task and the pose estimation of unseen object task together with [7, 17].
This evaluation is based directly on the raw pose results provided by the contenders and the official evaluation scripts of BOP challenges, both being publicly available on the Challenge website. Following the BOP Challenge, the evaluation is based exclusively on recall. Our method ranking is based on the mean of Recalls computed on MSPD and MSSD accuracies. As discussed in Section 2, we do not consider VSD metric as it is now omitted for new tasks in BOP.
@@ Results.
Table 1 reports the results of our evaluations. We can observe a large impact of the ground truth on the methods ranking since only the first place in the ranking remains unchanged. For example, regarding the ranking of the pose estimation task, the method Surfemb [6] moves from the 19th place with the original ground truth annotation to the 5th place with our annotation, gaining 14 places. For pose estimation of unseen objects, the difference of ranking can reach 8 places.
We observe large impacts on the recall performances as well. Whereas all the 20 first contenders of pose estimation task have a mean recall up to 80% and even up to 90% for 5 methods with the original annotations, only the 4 methods exceed this score with our annotation, with a majority of method having less than 65%. Similarly, all the 14 top contenders for the pose estimation of unseen objects task have a mean recall that exceed 50% with original annotation whereas only the best one exceed this score once evaluated with our annotations.
It is also interesting to note that the ranking order of variants 1.3, 1.4 and 1.5 of GDRNPP (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth. It underlines that inaccurate ground truth annotation can induce erroneous conclusions in ablation studies.
@@ 5.4 Pose Distribution Evaluation
In this experiment, we provide the first evaluation of 6D pose distribution estimation methods on a dataset of real images.
@@
Similarly to Section 5.3, the methods were trained on the train set of T-LESS that includes both photo-realistic synthetic images and real images.
The ground truth used for this dataset is the same than the one of Section 5.3 and is constituted of a discreet set of 6D pose obtained with the 6D pose annotation process introduced in Section 3. The performances of the methods are evaluated in Precision and Recall, using accuracy measures MPD and MSD as introduced in Section 4.
More details on methods processing and results formatting are given in the supplementary material.
@@ Results
We evaluated SpyroPose [7] and LiePose-Diffusion [17] since they provide 6D pose distribution, unlike [30, 16] that provide distribution only over the rotation.
Quantative results are reported on Table 2 whereas qualitative results are provided in the supplementary material.
It appears that LiePose-Diffusion outperforms SpyroPose. Without surprise, performances decrease when MSD accuracy is used since those two methods use RGB images only.
Methods $\mathbf{P_{MPD}}$ $\mathbf{R_{MPD}}$ $\mathbf{P_{MSD}}$ $\mathbf{R_{MSD}}$
SpyroPose [7] 69.9 69.9 55.2 65.9
LiePose-Diff [17] 75.4 94.1 48.1 85.3
Table 2: Comparisons of pose distribution estimation methods on T-LESS using our ground truth pose distributions. So far, these methods were evaluated quantitatively on synthetic images only and only qualitatively on real images.
@@ 6 Limitations
While our solution improves the pose annotations, the resulting pose distribution depends on some parameters (tolerance factors, sampling of the SE(3) space, etc.). We selected their values carefully. We let the reader check the visualizations in the supplementary material.
Our annotation method relies on a geometric analysis of the scene. It is possible that the sensor resolution, sensor noise, field of view, or motion blur may also affect the visibility of the disambiguating parts. Considering these effects would improve even more the ground truth.
@@ 7 Conclusion
For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, the latter being completed with a per-object symmetry pattern to transform this unique pose into a distribution. We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting distribution ground truth and in the resulting performances evaluations.
We then proposed a method to annotate 6D pose distribution with a per-image analysis of the object symmetries. We illustrated that the resulting ground truth is more accurate. Moreover, when using this ground truth to re-evaluate current state-of-the-art methods, we showed that the ranking of these methods changes drastically. We also introduced metrics to evaluate methods that estimate a pose distribution and provided the first evaluation of such methods on real data.
@@ 1.2 Differences between Original Annotations and Ours
Figure 4: Visualization of annotation changes compared to the T-LESS original annotations. Top: T-LESS objects with their identifier. Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages). T-LESS assumes full rotational symmetries, while our annotation method captures more complex symmetry patterns. Only Object 18 is perfectly symmetrical and our method retrieves the same poses as the original annotations. For the other objects, in particular the objects with complex symmetry patterns like the fist 4 objects. Our annotations significantly change the original annotations. The objects annotated as ’circular’ in BOP are highlighted in orange.

### v2

@@ 5.2 Qualitative validation of the method
RGB LiePose Diffusion [19] 60.1 92.2 76.1 1 48.3 76.1 62.2 -13.9 2( $\downarrow$-1)
SpyroPose [9] 61.2 75.3 68.3 2 57.8 71.6 64.7 -3.6 1( $\uparrow$1)
Table 1: Results of BOP 2023 on T-LESS [13] and YCB-V [47] with our annotations. We report the top contenders results of the pose estimation of unseen objects challenge and the pose estimation challenge evaluated with $\mathbf{MSSD}$ and $\mathbf{MSPD}$ on both the official ground truth (object-wise) and our ground truth (image-wise). We rank the methods with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ (similar to BOP, except we exclude $\mathbf{VSD}$ as BOP starts to abandon it) and re-rank them with the mean of $\mathbf{MSSD}$ and $\mathbf{MSPD}$ computed on our ground truth distributions. The drop in the scores highlighted by the $\mathbf{Loss}$ column ($\mathbf{Loss}=\mathbf{Mean}_{\mathbf{BOP-Distrib}}-\mathbf{Mean}_{\mathbf{BOP}}$) produces drastic changes in the rankings. We also show the same scores for the best mode of distribution estimation methods, however, both of them use ground truth bounding boxes instead of a detector, so direct comparison would be unfair. This highlights the importance to consider image level symmetries as we proposed, to evaluate accurately pose estimation methods. We also processed YCB-V. The $\mathbf{Loss}$ is much lower as there are fewer symmetrical objects. Names of the methods are automatically processed from BOP csv result files.
@@ 5.3 Single Pose Estimation Evaluation
Since May 2023, raw results of BOP Challenge [17] public submissions have become publicly available to allow in-depth analysis. In this experiment, we use them to compare the performance evaluations of state-of-the-art methods on T-LESS for both the original 6D annotations and our annotations. We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks. We also included two methods for the multi-modal pose distribution estimation task [9, 19] for which we kept only the mode with highest confidence. We consider their full predicted distributions in the next subsection. Since these methods are not part of the BOP challenge, we retrained them on the T-LESS train set.
@@
We evaluate the current (July 2024) top contenders of the BOP Challenge on the T-LESS dataset [13] on the pose estimation task and the pose estimation of unseen object task together with [9, 19]. We conduct a similar evaluation with YCB-V [47].
This evaluation is based directly on the raw pose results provided by the contenders and the official evaluation scripts of BOP challenges, both being publicly available on the Challenge website. Following the BOP Challenge, the evaluation is based exclusively on recall. Our method ranking is based on the mean of Recalls computed on MSPD and MSSD accuracies.
@@ Results.
Table 1 reports the results. We can observe a large impact of the ground truth on the methods results (see Loss column) and ranking. In terms of ranking, the well established pose estimation task face the biggest changes, with gdrnppdet-pbrrealmegapose-multihyp moving from the 12th to the 3rd position for RGB methods and surfemb-pbr-rgbd-lin [8] moving from the 16th to the 5th place for the RGB-D methods. For pose estimation of unseen objects, the difference of ranking can reach 6 places.
We observe large impacts on the recall performances as well. Whereas 29 of the 30 contenders of pose estimation task have a mean recall up to 80% and even up to 90% for 6 methods with the original annotations, only 4 methods exceed this score with our annotations, with a majority of methods having less than 65%. Similarly, most of the contenders for the pose estimation of unseen objects task have a mean recall that exceeds 50% with the original annotations whereas only the 2 best RGB-D methods exceed this score once evaluated with our annotations. As shown in Figure 1 (Case 3), the image-wise annotation rejects poses when the image is not ambiguous. The Loss is explained by inaccurate pose estimates that were mistakenly validated by previous object-wise annotations (see supplementary material for visualizations).
It is also interesting to note that the ranking order of variants v1.2, v1.3 and v1.4 of GDRNPP RGB-D (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth. It underlines that inaccurate ground truth annotation can induce erroneous conclusions in ablation studies.
@@ 5.4 Pose Distribution Evaluation
We provide now the first evaluation of 6D pose distribution estimation methods on a dataset of real images.
@@
Similarly to Section 5.3, the methods were trained on the train set of T-LESS. The ground truth used for this dataset is the same than the one of Section 5.3 and is constituted of a discrete set of 6D pose obtained with the 6D pose annotation process introduced in Section 3. The performances of the methods are evaluated in Precision and Recall, using accuracy measures MPD and MSD as introduced in Section 4. More details on methods processing and results formatting are given in the supplementary material.
@@ Results
We evaluated SpyroPose [9] and LiePose-Diffusion [19] (either PBR + real images and PBR only to compare with SpyroPose) since they provide 6D pose distributions, unlike [33, 18] that provide distributions only over the rotation. To show genericity, we also evaluated the top 3 RGB methods of single pose estimation task (with image-wise annotations). Their distribution is considered as a Dirac distribution on the single estimate. Results are given on the complete dataset, and the subset of objects with cylindrical symmetry (1-4, 13-18, 24 and 30, as illustrated in supplementary material).
Quantitative results are reported on Table 2 whereas qualitative results are provided in the supplementary material. It appears that LiePose-Diffusion outperforms SpyroPose, except for the Precision MSD. Lower accuracy are obtained with MSD error since those methods use RGB images only and no depth. Single pose methods outperform distribution methods in terms of Precision, as they are optimized for that. Yet, distribution methods produce much stronger Recalls, even more on only symmetrical objects, highlighting their ability to retrieve multiple meaningfull poses from the targeted image-wise distribution.
Methods Train $\mathbf{P_{MSD}}$ $\mathbf{R_{MSD}}$ $\mathbf{P_{MPD}}$ $\mathbf{R_{MPD}}$
T-LESS [13] Complete dataset
Distrib SpyroPose [9] PBR 32.8 48.1 55.9 55.5
LiePose-Diffusion [19] PBR 24.6 71.4 61.2 89.5
LiePose-Diffusion [19] PBR + real 29.9 75.4 68.4 90.6
leroy-fuseocclu-rgb PBR + real 55.4 35.7 80.9 54.2
gdrnppdet-pbrrealgenflow-multihypo-rgb [44, 28] PBR + real 52.4 34.9 77.7 52.2
Single cosypose-eccv20-syntreal-8views [23] PBR + real 43.5 25.7 53.1 30.9
T-LESS [13] Only symmetrical objects
Distrib SpyroPose [9] PBR 31.5 42.7 70.5 60.8
LiePose-Diffusion [19] PBR 18.3 61.6 63.6 87.3
LiePose-Diffusion [19] PBR + real 24.1 67.0 70.3 88.1
leroy-fuseocclu-rgb PBR + real 52.4 25.7 86.1 45.3
gdrnppdet-pbrrealgenflow-multihypo-rgb [44, 28] PBR + real 45.8 20.8 76.5 37.8
Single cosypose-eccv20-syntreal-8views [23] PBR + real 29.1 6.5 40.0 10.5
Table 2: Comparison of pose distribution estimation methods on T-LESS using our ground truth pose distributions.
@@ 6 Limitations
While our solution improves the pose annotations, the resulting pose distribution depends on some parameters (tolerance factors, sampling of the SE(3) space, etc.). We selected their values carefully. We invite the reader to check the visualizations in the supplementary material.
Our annotation method relies on a geometric analysis of the scene. It is possible that the sensor resolution, sensor noise, field of view, or motion blur may also affect the visibility of the disambiguating parts. Considering these effects would improve the ground truth even further.
@@ 7 Conclusion
For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, completed by a per-object symmetry pattern to transform this unique pose into a distribution. We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting distribution ground truth and in the resulting performances evaluations.
We then proposed a method to annotate 6D pose distribution with a per-image analysis of the object symmetries. We illustrated that the resulting ground truth is more accurate. Moreover, when using this ground truth to re-evaluate current state-of-the-art methods, we showed that the ranking of these methods changes drastically. We also introduced metrics to evaluate methods that estimate a pose distribution and provided their first evaluation on real data.
@@ 1.2 Differences between Original Annotations and Ours
Figure 4: Visualization of annotation changes compared to the T-LESS original annotations. Top: T-LESS objects with their identifier. Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages). T-LESS assumes full rotational symmetries, while our annotation method captures more complex symmetry patterns. Only Object 18 is perfectly symmetrical and our method retrieves the same poses as the original annotations. For the other objects, in particular the objects with complex symmetry patterns like the first 4 objects. Our annotations significantly change the original annotations. The objects annotated as ’circular’ in BOP are highlighted in orange.